Suppose you know that a company’s stock currently sells for $53 per share and the required return on the stock is 10 percent. You also know that the total return on the stock is evenly divided between a capital gains yield and a dividend yield. It's the company’s policy to maintain a constant growth rate in its dividends.

What is the current dividend per share? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Current Dividend per share = __________
Answer #1

We know the stock has a required return of 10 percent, and the dividend and capital gains yield are equal, so:

Dividend yield = 1/2(.10) = .05 = Capital gains yield

Now we know both the dividend yield and capital gains yield. The dividend is simply the stock price times the dividend yield, so:

D1 = .05($53)

D1 = $2.65

This is the dividend next year. The question asks for the dividend this year. Using the relationship between the dividend this year and the dividend next year:

D1 = D0(1 + g)

$2.65 = D0(1 + .05)

D0 = $2.65 / 1.05

D0 = $2.52